# Exploratory Causal Inference in SAEnce

### v1

@@ E.2 In–depth analysis: full semi–synthetic results
    Unknown number of effects ($r$). Each method returns its own set of significant codes at level $\alpha$ or simply the Top-K. We then report Precision, Recall, and IoU against the ground–truth affected codes (Section D.1).
@@
    Known number of effects ($r$). We assume to know the true number of effects, and we just look at $r$- highest effect among each method. We again compute Precision, Recall, and IoU (namely, we apply Top-2 selection on top of other methods).
As detailed in Appendix D.1, we vary (i) the sample size $n\in\{30,50,100,250,500,1000\}$ and (ii) the ATE magnitude $\tau\in\{0.1,\dots,0.8\}$, holding the semi–synthetic DGP and SAE training protocol fixed. Each cell aggregates 10 random seeds (RCT re–draws and SAE initializations).
Main takeaways.
Across both regimes and over the entire grid, NES maintains high Precision and IoU while matching the best Recall of baselines. When the experiment power increases (larger $n$ or $\tau$), vanilla tt–tests and classical multiplicity corrections (FDR/Bonferroni) exhibit the significance–collapse behavior: Recall saturates but Precision drops sharply as leakage neurons become significant, driving IoU toward zero. Enforcing the correct cardinality ($r$ known) mitigates over–selection but does not resolve entanglement: baselines still replace a true effect with a leakage surrogate in later picks, keeping Precision <0.5<\!0.5 in the high–power regime. In contrast, NES’s residual stratification peels one principal effect component per round and then stops, preserving interpretability.
@@
E.3 Ablation I: No Causal Effect
We repeat the semi-synthetic evaluation of Section E.2 but set the true ATE to zero, namely $\tau=0$ factors. In this regime, a well-calibrated discovery procedure should return no significant neurons.
@@ E.3 Ablation I: No Causal Effect
Figure 10: Zero-effect ablation . Number of discovered neurons by method when ATE is 0.
@@
We keep the data-generating process, foundation model, SAE training, and testing grid over sample sizes $n$ identical to Section  E.2, changing only the interventional contrast to $\mathrm{ATE}=0$. For each method, we record the number of discoveries per run. Across all sample sizes, NES returns an empty set: in the first iteration, no neuron survives Bonferroni at level $\alpha/m$, and the recursion halts. Furthermore, both Bonferroni and FDR also yield essentially zero discoveries. In contrast, the uncorrected $t$-test produces spurious positives (false discoveries), and Top-$k$ necessarily reports $k$ indices by design, labeling pure noise as significant. This behavior matches our theoretical intuition: with $\tau=0$ there is no effect vector to leak into entangled coordinates, so the paradox of Sec. 3 does not arise; procedures that control multiplicity (NES via its first-step Bonferroni gate, Bonferroni, and FDR) appropriately abstain, whereas selection rules that ignore multiplicity (Top-$k$, plain $t$-tests) over-discover.
E.4 Ablation II: Testing in NES
We compare three per-round gates inside NES (Alg. 1): Bonferroni, FDR, and $t$-test. Same setup as Sec. E.2; only the multiplicity rule changes while recursion and residual stratification are unchanged. NES-Bonf. delivers the cleanest recoveries: highest precision/IoU and exact stopping at $r$ effects when powered; under $\mathrm{ATE}{=}0$ it returns none (cf. Ablation E.3). NES-$t$ is most exploratory for small sample size and effect magnitude but over-selects as power grows, i.e., Paradox of Exploratory Causal Inference.
Recommendation. Prefer a multi-hypothesis testing correction, i.e., Bonferroni/FDR, when the power of the experiment is high, while consider $t$-test for a more explorative approach in low power regime.
Figure 11: Testing in NES. Bonferroni: best precision/IoU and exact stopping; FDR: higher sensitivity in low power, minor over-selection; $t$-test: exploratory but prone to over-discovery as power increases.
E.5 Ablation III: AIPW vs. Associational Difference
Throughout the paper, our per-neuron hypothesis test uses the associational difference (AD), i.e., a two-sample $t$-test on the treated–control difference in means. In randomized trials, AD is unbiased for the ATE, but it is not semiparametrically efficient. A standard variance–reduction alternative is Augmented Inverse Propensity Weighting (AIPW; Robins et al., 1994), which orthogonalizes the estimator against misspecification of either the propensity score or the outcome regression.
Setup.
For each code $j$, let $Z_{ij}$ be its activation for unit $i$, $T_{i}\!\in\!\{0,1\}$ the treatment, and $W_{i}$ observed exogenous causes. We compute the AIPW pseudo-outcome
where $\pi(W)\!=\!\Pr(T\!=\!1\mid W)$ (known and constant $\pi=0.5$ in our RCT), and $\hat{\mu}_{tj}(W)\approx\mathbb{E}[Z_{j}\mid T\!=\!t,W]$ is a nuisance regression. The AIPW estimate of the code-level ATE is $\hat{\tau}^{\text{AIPW}}_{j}=\frac{1}{n}\sum_{i}\tilde{Z}_{ij}$; we test $H_{0}:\tau_{j}=0$ via a one-sample $t$-test on $\{\tilde{Z}_{ij}\}_{i}$ with robust variance.
Results.
Figure 12 compares AD vs. AIPW on the semi-synthetic benchmark across sample size $n$ and effect magnitude $\tau$. In our setting—with a truly randomized treatment ($\pi=0.5$) and a single binary covariate WW—AIPW yields only marginal efficiency gains: Precision/Recall/IoU curves are essentially overlapping, with small stability improvements for AIPW at the smallest $n$. Crucially, orthogonalization affects variance but does not resolve entanglement: the significance–collapse phenomenon for standard multi-testing (Section 3) persists under AIPW, and NES retains its advantage because its benefit comes from recursive stratification (disentangling residual effects), not from how the first-step mean contrast is estimated.
Takeaways.
(i) In pure RCTs with weak, low-dimensional $W$, AD is competitive and simpler. (ii) AIPW can be preferred when richer exogenous information is available (higher-dimensional $W$, imbalance, or mild protocol deviations), where its variance reduction can translate into earlier detection of the leading effect; (iii) regardless of AD or AIPW, NES’s stratified recursion is the key to avoiding over-discovery under entanglement.
Figure 12: AIPW vs. AD on semi-synthetic RCTs. Precision, Recall, and IoU when replacing the per-neuron associational difference (AD) with AIPW (Eq. 36) for baselines and the first NES step.

### v2

@@ E.2.3 Varying Structural Causal Model
    Unknown number of effects ($r$). Each method returns its own set of significant codes. We then report Precision, Recall, and IoU against the ground–truth affected codes.
@@
    Known number of effects ($r$). We assume to know the true number of effects, and we just look at the $r$ highest effects for each method. We again compute Precision, Recall, and IoU (namely, we apply top-$2$ selection on top of other methods).
As detailed in Appendix D.1, we vary (i) the sample size $n\in\{30,50,100,250,500,1000\}$ and (ii) the ATE magnitude $\tau\in\{0.1,0.2,\dots,0.8\}$, holding the semi–synthetic DGP and SAE training protocol fixed. Across both regimes and over the entire grid, NES maintains high Precision and IoU while matching the best Recall of baselines. When the experiment power increases (larger $n$ or $\tau$), vanilla tt–tests and classical multiplicity corrections (FDR/Bonferroni) exhibit the significance–collapse behavior: Recall saturates but Precision drops sharply as leakage neurons become significant, driving IoU toward zero. Enforcing the correct cardinality ($r$ known) mitigates over–selection but does not resolve entanglement: baselines still replace a true effect with a leakage surrogate in later picks, keeping Precision <0.5<\!0.5 in the high–power regime. In contrast, NES’s residual stratification peels one principal effect component per round and then stops, preserving interpretability.
@@
No effect
We repeat the experiments described in Appendix D, removing both the effects, namely imposing $\tau=0$. In this setting, a valid discovery procedure should return no significant neurons.
@@
Figure 32: Zero-effect ablation. Number of significant neurons retrieved by method when no effects present, i.e., optimal if not retrieving anything.
@@
We keep the data-generating process, foundation model, SAE training, and testing grid over sample sizes $n$ as described in Appendix D, changing only the interventional contrast to $\mathrm{ATE}=0$. For each method, we record the number of discoveries per run. Across all sample sizes, NES consistently returns an empty set: in the first iteration, no neuron survives Bonferroni at level $\alpha/m$, and the recursion halts. Furthermore, both Bonferroni and FDR also yield essentially zero discoveries. In contrast, the uncorrected $t$-test produces spurious positives (false discoveries), and top-$k$ necessarily reports $k$ indices by design, labeling pure noise as significant. This behavior matches our theoretical intuition: with $\tau=0$ there is no effect vector to leak into entangled coordinates, so the paradox discussed in Section  3 does not arise; still procedures that control multiplicity (NES via its first-step Bonferroni gate, Bonferroni, and FDR) appropriately abstain, whereas selection rules that ignore multiplicity (top-$k$, vanilla $t$-tests) over-discover, collapsing the method precision.
Opposite effects
To stress–test effect identification under more complex causal structure, we extend the main RCT simulation by allowing the two outcome components $Y_{1}$ and $Y_{2}$ to have different treatment effects (even complementary). As before, each synthetic unit corresponds to a real CelebA image whose attributes match the realized $(Y_{1},Y_{2},W)$ triple. We fix $T\!\sim\!\mathrm{Bernoulli}(0.5)$ and $W\!\sim\!\mathrm{Bernoulli}(0.5)$, and we retain the co–effect model described in Appendix D, but now vary the two ATEs over a grid of ordered pairs $(\tau_{1},\tau_{2})\in\{(0.2,0.8),(0.3,0.7),\dots,(0.8,0.2)\}$, thus controlling the gap between the two causal effects. For each pair, we set the treatment–induced shift for $Y_{2}$ as
$p^{(Y_{2})}_{1}=0.5+\tfrac{\tau_{2}}{2},\qquad p^{(Y_{2})}_{0}=0.5-\tfrac{\tau_{2}}{2},$ (37)
while $Y_{1}$ additionally depends on $W$:
$\Pr(Y_{1}{=}1\mid T{=}t,W{=}w)=\begin{cases}p^{(Y_{1})}_{11}=0.5+\tfrac{\tau_{1}}{2},&(t{=}1,w{=}1),\\[2.0pt] p^{(Y_{1})}_{10}=0.2+\tau_{1},&(t{=}1,w{=}0),\\[2.0pt] p^{(Y_{1})}_{01}=0.5-\tfrac{\tau_{1}}{2},&(t{=}0,w{=}1),\\[2.0pt] p^{(Y_{1})}_{00}=0.2,&(t{=}0,w{=}0).\end{cases}$ (38)
This yields a controlled spectrum of settings in which $Y_{1}$ and $Y_{2}$ express treatment effects of varying and potentially opposite magnitudes. For each $(\tau_{1},\tau_{2})$ and sample size $N\!\in\!\{500,5000\}$, we draw $(T,W,Y_{1},Y_{2})$, attach the matching real image, and apply NES and baselines to evaluate precision, recall, and IoU across the full two–effect grid. The results are reported in Figure 33-34, and consistently replicate the evidence discussed in the main experiments, i.e., NES overcoming the Paradox of Exploratory Causal Inference, as opposed to any other baselines. As expected, with enough experiment power, e.g., N=5  [^0] N=5\ 000, all the methods consistently return the same results for any effects couple, while with lower power, e.g., $N=500$, both NES and some vanilla baseline can struggle in retrieving all the the effects (perfect recall) if one or both the signals are very limited (as for the main experiments).
Figure 33: Ablation opposite effects ($N=500$). Precision, Recall, and IoU in effect identification varying the two effects difference in the main structural causal model (fixing $N=500$). NES consistently replicates the main results, still overcoming the paradox of Exploratory Causal Inference.
Figure 34: Ablation opposite effects (N=5  [^0] N=5\ 000). Precision, Recall, and IoU in effect identification varying the two effects difference in the main structural causal model (fixing N=5  [^0] N=5\ 000). NES consistently replicates the main results, still overcoming the paradox of Exploratory Causal Inference.
Varying propensity score
Finally, we repeat the main experiments (fixing treatment effect $\tau=0.6$ and sample size $N\in\{500,5\ 000\}$) varying the treatment assignment probability $\mathbb{P}(T=1)\in\{0.2,0.4,0.6,0.8\}$, i.e., propensity score with no dependence on the covariates being a Randomized Controlled Trial. We report the results in Figures 35-35. Once again, both NES and the other baselines replicate the findings described for the main experiments in Section 6. Particularly, NES consistently overcomes the Paradox of Exploratory Causal Inference, which is not the case for the other baselines. The results are consistent and almost invariant with respect to $\mathbb{P}(T=1)$, even if, as expected, edge values of the propensity score, e.g., 0.2 and 0.8, can slightly affect the method’s efficiency.
Figure 35: Ablation unbalanced treatment assignment ($N=500$, $\tau=0.6$). Precision, Recall, and IoU in effect identification varying the treatment assignment probability with sample size $N=500$. NES consistently replicates the main results in all the regimes, still overcoming the paradox of Exploratory Causal Inference.
Figure 36: Ablation unbalanced treatment assignment (N=5  [^0] N=5\ 000, $\tau=0.6$). Precision, Recall, and IoU in effect identification varying the treatment assignment probability with sample size N=5  [^0] N=5\ 000. NES consistently replicates the main results in all the regimes, still overcoming the paradox of Exploratory Causal Inference.
E.3 Additional Baselines
Finally, we further include here two additional non-trivial, still heuristic baselines to compare our method with. Particularly inspired by Causal Feature Learning [chalupka2017causal], which attempts to abstract effects by clustering with respect to $\mathbb{P}(T\mid X=x)$, we propose to consider an additional neural effect selector obtained by regressing the treatment $T$ on the SAE representation of $X$, and filtering only the relevant codes. The idea is that the (by design) non-visible treatment information is indirectly retrievable from the effect measurement, and that such information flows through the anti-causal path $Y\leftarrow T$.
LASSO.
We consider a sparsified logistic regression model (recall that $T$ is binary) with an $\ell_{1}$ penalty to enforce sparsity on the SAE coefficients. Given SAE codes $Z\in\mathbb{R}^{N\times m}$ and treatment labels $T\in\{0,1\}^{N}$, logistic regression models
$p(T_{i}=1\mid Z_{i},\beta)=\sigma\!\left(\beta_{0}+Z_{i}^{\top}\beta\right)=\frac{1}{1+\exp\!\left(-\beta_{0}-\sum_{j=1}^{m}Z_{ij}\beta_{j}\right)},$ (39)
and fits parameters by minimizing the penalized negative log-likelihood
$\hat{\beta}=\arg\min_{\beta\in\mathbb{R}^{m}}\Bigg\{-\sum_{i=1}^{N}\Big[T_{i}\log p_{i}+(1-T_{i})\log(1-p_{i})\Big]+\lambda\lVert\beta\rVert_{1}\Bigg\},$ (40)
where $\lambda>0$ controls sparsity. The resulting selected neurons are
$\texttt{S}_{\text{lasso}}=\{j\;:\;\hat{\beta}_{j}\neq 0\}.$ (41)
Stability-Selected LASSO.
To mitigate instability due to correlated SAE codes, we additionally consider stability selection. Given a fixed regularization strength $\lambda$ (e.g., tuned via cross-validation), we repeatedly refit the LASSO model on bootstrapped subsamples of the data. For each bootstrap iteration $b=1,\dots,B$, let $\hat{\beta}^{(b)}$ be the fitted coefficients. We then define a stability score
$\Pi_{j}=\frac{1}{B}\sum_{b=1}^{B}\mathbb{I}\!\left(\hat{\beta}^{(b)}_{j}\neq 0\right),$ (42)
and select those neurons consistently appearing across many subsamples:
$S_{\text{stable}}=\{j\;:\;\Pi_{j}\geq s\},$ (43)
where $s\in(0,1)$ is a high-confidence threshold. In practice, we consider $s=0.6$.
We replicate once again the main experimental setup described in Section D, additionally including the two novel baselines, and report the results in Figure 37. Interestingly, the new baselines reach comparable performance to vanilla multiple-testing methods, despite explicitly exploiting the anti-causal path and never measuring the neural effect directly. With sufficient experimental power, they also reach full-effect recall, but—similarly to vanilla multi-testing—still suffer from the Paradox of Exploratory Causal Inference. Indeed, we are not aware of any theoretical result guaranteeing that these heuristics should recover the standalone ground-truth affected outcomes among all candidate neurons. Stability selection appears to boost vanilla LASSO performance, and with sufficient power, these approaches slightly outperform vanilla multi-testing (although precision remains unsatisfactory, i.e., $\ll 1$).
Figure 37: Additional Baselines from Causal Feature Learning. Precision, Recall, and IoU in effect identification, including two baselines inspired by effect abstraction in Causal Feature Learning using LASSO with or without stability selection. NES consistently replicates the main results in all the regimes, still overcoming the paradox of Exploratory Causal Inference.